# Nonlinear phase synchronization and the role of spacing in shell models

### v1

@@ VI Conclusion
A detailed analysis of the dynamics of triadic phases in a number of shell models, interpreting them as chains of triads, where each triad is a nonlinear oscillator, reveals how synchronizability plays a key role in the statistics of the energy cascade. In particular, our result show that phase synchronized events are associated with bursts of energy transfer, which determine the intermittency of the system. Our investigation, which is in line with previous results on Burgers equation that links intermittency to phase dynamics, provides new evidence that stronger intermittency is related to increased triadic phase coherence, as evidenced in a shell model of three dimensional Navier Stokes equations.
@@
In the classical GOY model, with fixed inter-shell spacing, $g=2$, the system exhibits the “usual level” of intermittency. Decreasing inter-shell spacing, while keeping the same physical quantities conserved , we systematically examine how each scaling exponent changes continuously, ultimately approaching a Burgers-equation like dynamical scaling in the continuum limit. This scenario allows us to investigate the casual link between synchronization and intermittency, demonstrated by the use of a Kuramoto order parameter to quantify phase coherence among the triad population of a shell model. The PDFs of the order parameter $R$ reveal an increased likelihood for fully coherent events as $g$ decreases, accompanied by a loss of scale universality in the statistics of triadic phases. Complementary to this, the average phase (i.e. the phase associated with the complex Kuramoto parameter) distribution becomes more peaked around $\pi/2$, which is incidentally the value that maximizes forward energy transfer.
@@
In shell models, extreme energy flux events are associated with coherent pulses, propagating across the inertial range. We observe that, as $g$ decreases, those pulses becomes more localized and significantly more intense but less frequent. In order to study the correlation between the dynamics of such burst-like events and the coherence of the triad chain, a novel order parameter was introduced, to quantify local phase coherence. This weighted Kuramoto parameter, derived from the right hand side of the triadic phase evolution equation, captures how a coherent pulse acts as a synchronization domino effect on consecutive triadic phases. Specifically, the passage of a coherent event is preceded by a rapid increase in the local coherence parameter that synchronizes a cluster of triads in a partially phase locked state, i.e. $R_{n}\simeq 1$, or $\varphi_{n}=\pi/2,$ that maximizes the energy transfer between aligned triads, suggesting a casual relation between the local synchronization and the coherent burst in energy transfer. The strong correlation between this phase locked state and large deviations in energy flux, can also be evidenced by their joint statistics.
@@
The PDFs of the local Kuramoto parameter also shows that local coherence precedes global synchronization. As intermittency increases, the coherent events generate longer-lasting phase-locked states, as demonstrated by the elevated probability of observing $R_{n}\simeq 1$. This suggests that accumulation of energy at large scales due to an incoherent phase distribution at small scales, triggers the instanton, whose forced passage reorganizes the phases in such a way that the energy can continue to flow along this opened “path” for a while, which is then slowly de-synchronized from small scales, and as the desynchronization propagates to large scales and becomes significant, the “path” gets closed again, which in turn causes the accumulation of energy at large scales, eventually repeating the whole cycle.
@@
As these observations show, the local Kuramoto parameter has proven to be a useful proxy for detecting coherent events. It may be possible to generalize this for detecting paths of burst events in models with spatial structure such as binary trees or more complex $k$-space structure such as nested polyhedra models, and maybe eventually direct numerical simulations. However, it should be mentioned that defining a local phase coherence indicator in nonlinear systems involving many interconnected triads is a nontrivial task.
@@
As a minimal example of the case with interconnected triads, we investigated triadic phase dynamics in helical shell models constructed from two sequences of helical variables, allowing for different classes of chiral interactions, focusing in particular on the SM3 class, which results in two intertwined chains that are topologically non-separable. We observe that contrary to the single chain case, here decreasing the intershell spacing leads to a reduction of intermittency, with the system exhibiting self similar mono-fractal scaling as it approaches the continuum limit. However, as before we find that the phase coherence correlates with intermittency, as shown by the order parameters, even though the dependence of that with $g$ is the inverse of what happens in the GOY case.
@@
Finally, we have addressed the role of triadic phases on the inverse energy cascade, considering two specific cases, represented by a non-local helical shell model and the model investigated in Ref. [49]. Our results show that inverse cascades are characterized by a PDF of triadic phases, peaked around $\varphi_{n}=-\pi/2$ , which maximizes the inverse energy transfer. However the order parameters show that the inverse cascades (at least the ones that we considered), are non intermittent, and present a significantly lower levels of phase coherence in contrast to the forward cascades. This is also in agreement with the emerging picture that phase dynamics and intermittency in reduced models are intimately connected.
@@
In short, our findings provide new evidence that phase synchronization acts as a driver of intermittency and extreme events in shell models. We introduced novel parameters to quantify the local phase coherence in these models, which can be extended to similar cascade models using a network perspective.
@@ ACKNOWLEDGMENT
This work has benefited from a grant managed by the Agence Nationale de la Recherche (ANR), as part of the program ‘Investissements d’Avenir’ under the reference (ANR-18-EURE-0014) and has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion) and within the framework of the French Research Federation for Fusion Studies.
@@ Appendix A Numerical details
A special note concerns the choice of the forcing term. The common approach in the literature is to apply either a constant or random forcing. In the specific case discussed in Sec. III to investigate the role of phase self-organization, particularly in relation to phase-syncronized events approaching the continuum limit, we implemented a forcing term that does not inject any phase into the system. This is achieved by ensuring that the term $Im\left(\frac{f_{n}}{\rho_{n}}e^{-i\theta_{n}}\right)$ in the right-hand side of Eq.4 is identically zero, thus we fixed $f_{n}=Ae^{i\theta_{n}}$, with $A=0.5$ , in all the runs corresponding to Table 1. Another possible choice,$f_{n}=\epsilon/u_{n}^{*}$, which does not inject any phase and fix a constant energy injection $\epsilon$, as discussed in Ref. [57] in order to study the statistics of inter-event times. We note that for the result of present analysis there are no significant differences among these two choices.
@@
$N$ 164 113 88 62 50 42 38 36 33 31 29
$\nu$ $5\times 10^{-9}$ $4\times 10^{-9}$ $8\times 10^{-9}$ $14\times 10^{-9}$ $9\times 10^{-9}$ $8\times 10^{-9}$ $5\times 10^{-9}$ $9\times 10^{-10}$ $7\times 10^{-10}$ $5.5\times 10^{-10}$ $5.5\times 10^{-10}$
@@
Table 2: Numerical parameters corresponding to the different runs of the local Helical shell models.$f_{3}=(1+1j)\times 7\times 10^{-1}$

### v2

@@ VI Conclusion
A detailed analysis of the dynamics of triadic phases in several shell models, interpreted as chains of interacting triads, where each triad is considered as a nonlinear oscillator, reveals how synchronizability plays a key role in the statistical properties of the energy cascade. In particular, our results show that phase-synchronized events are associated with bursts of energy transfer, which determine the intermittency of the system. Our investigation, which is in line with previous results on Burgers equation that links intermittency to phase dynamics, provides new evidence that stronger intermittency is related to enhanced triadic phase coherence, which is demonstrated first in a shell model representation of the three dimensional Navier Stokes equations.
@@
In the classical GOY model, with fixed inter-shell spacing, $g=2$, the system exhibits the “usual level” of intermittency. Decreasing inter-shell spacing below a critical value, while keeping the same physical quantities conserved, we systematically examine how each scaling exponent changes continuously, ultimately approaching a Burgers-like dynamical scaling in the continuum limit. This transition is accompanied by a loss of scale invariance in the statistics of triadic phases and in the Kolmogorov multipliers. Remarkably, the enhancement of intermittency correction is associated with a breaking of the statistical symmetry of both quantities: while their PDFs are scale-independent at the “usual level” of intermittency, they lose this invariance as the intermittency increases, providing an interesting challenge to hidden-symmetry theories [69, 70].
This progressive enhancement of intermittency, controlled by the inter-shell spacing, allows us to investigate the connection between synchronization and intermittency, systematically using a Kuramoto order parameter to quantify global phase coherence among the triads of the shell model. The PDFs of the order parameter $R$ reveal an increased likelihood for fully coherent events as $g$ decreases, complementary to this, the average phase $\Phi$ (i.e. the phase associated with the complex Kuramoto parameter) distribution becomes more peaked around $\pi/2$, which is incidentally the value that maximizes forward energy transfer.
@@
In shell models, extreme energy flux events are associated with coherent pulses, propagating across the inertial range. We observe that, as $g$ decreases, those pulses becomes more localized and significantly more intense but less frequent. In order to study the correlation between the dynamics of such burst-like events and the phase coherence of the triad chain, a novel order parameter was introduced, to quantify local phase coherence. This weighted Kuramoto parameter, derived from the right-hand side of the triadic phase evolution equation, captures how a coherent pulse acts as a synchronization domino effect on consecutive triadic phases. Specifically, the passage of a coherent event is preceded by a rapid increase in the local coherence parameter that synchronizes a cluster of triads in a partially phase locked state, i.e. $R_{n}\simeq 1$, or $\Phi_{n}\simeq\pi/2,$ that maximizes the energy transfer between aligned triads, suggesting that the local synchronization is accompanied by the coherent burst in energy transfer. The strong correlation between this phase locked state and large deviations in energy flux, can also be evidenced by their joint statistics.
@@
The PDFs of the local Kuramoto parameter also show that local coherence precedes global synchronization. As intermittency increases, the coherent events generate longer-lasting phase-locked states, as demonstrated by the elevated probability of observing $R_{n}\simeq 1$. This suggests that accumulation of energy at large scales due to an incoherent phase distribution at small scales, triggers the instanton, whose forced passage reorganizes the phases in such a way that the energy can continue to flow along this opened “path” for a while, which is then slowly desynchronized from small scales, and as the desynchronization propagates to large scales and becomes significant, the “path” gets closed again, which in turn causes the accumulation of energy at large scales, eventually repeating the whole cycle.
@@
As these observations show, the local Kuramoto parameter can be a useful proxy for detecting coherent events that drive the energy cascade across range of scales. It may be possible to generalize this for detecting paths of burst events to models with spatial structure such as binary trees or more complex $k$-space structure such as nested polyhedra models, and maybe eventually to direct numerical simulations. In particular, the analysis of tradic phase dynamics in a 3D DNS could follow two different approaches: focusing on a single “path” in Fourier space, represented by a self-similar chain of connected triads in the spirit of spiral-chain models, or filtering DNS data into a set of shells and studying the triadic phase dynamics of those. However, it should be mentioned that defining a local phase coherence indicator in nonlinear systems involving many interconnected triads is a nontrivial task.
@@
As a minimal example of the case with interconnected triads, we investigated triadic phase dynamics in helical shell models constructed from two sequences of helical variables, allowing for different classes of chiral interactions, focusing in particular on the SM3 class, which consist of two intertwined chains that are topologically non-separable. We observe that, contrary to the single chain case, here, decreasing the inter-shell spacing leads to a reduction of intermittency, with the system exhibiting self-similar monofractal scaling as it approaches the continuum limit. However, as before we find that the phase coherence correlates with intermittency, as shown by the order parameters, even though the dependence of that with $g$ is the inverse of what happens in the GOY case.
@@
Finally, we have addressed the role of triadic phases in the inverse energy cascade, considering two specific cases, represented by a non-local helical shell model and the model investigated in Ref. 50. Our results show that inverse cascades are characterized by a PDF of triadic phases, peaked around $\varphi_{n}=-\pi/2$ , which maximizes the inverse energy transfer. However the order parameters show that the inverse cascades (at least the ones that we considered), are non-intermittent, and present significantly lower levels of phase coherence in contrast to the forward cascades. This is also in agreement with the emerging picture that phase dynamics and intermittency in cascade models are intimately connected.
@@
In short, our findings provide new evidence that phase synchronization acts as a driver of intermittency and extreme events in shell models. We introduced novel parameters to quantify the local phase coherence in these models, which can be extended to similar cascade models using a network perspective, and eventually to direct numerical simulations.
@@ ACKNOWLEDGMENT
This work has benefited from a grant managed by the Agence Nationale de la Recherche (ANR), as part of the program ‘Investissements d’Avenir’ under the reference (ANR-18-EURE-0014) and has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion) and within the framework of the French Research Federation for Fusion Studies. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, where part of the work on this paper was undertaken. This work was supported by EPSRC grant EP/R014604/1. In particular, the authors would like to thank Dr. Santiago J. Benavides for stimulating discussion on phase dynamics during the program.
DATA AVAILABILITY
The data that support the findings of this article are openly available at [71].
@@ Appendix A Numerical details
A special note concerns the choice of the forcing term. The common approach in the literature is to apply either a constant or random forcing. In the specific case discussed in Sec. III to investigate the role of phase self-organization, particularly in relation to phase-syncronized events approaching the continuum limit, we implemented a forcing term that does not inject any phase into the system. This is achieved by ensuring that the term $Im\left(\frac{f_{n}}{\rho_{n}}e^{-i\theta_{n}}\right)$ on the right-hand side of Eq. (4) is identically zero, thus we fixed $f_{n}=Ae^{i\theta_{n}}$, with $A=0.5$ , in all the runs corresponding to Table 1. Another possible choice, $f_{n}=\epsilon/u_{n}^{*}$, which does not inject any phase and fix a constant energy injection $\epsilon$, as discussed in Ref. 58 in order to study the statistics of inter-event times. We note that for the result of present analysis there are no significant differences among these two choices.
@@
NN 164 113 88 62 50 42 38 36 33 31 29
$\nu$ $5\times 10^{-9}$ $4\times 10^{-9}$ $8\times 10^{-9}$ $14\times 10^{-9}$ $9\times 10^{-9}$ $8\times 10^{-9}$ $5\times 10^{-9}$ $9\times 10^{-10}$ $7\times 10^{-10}$ $5\times 10^{-10} 5.5$$5\times 10^{-10}$
@@
[…]